# Gradual smoothing: strong hypercontractivity and logarithmic Sobolev inequalities

### v1

@@ 6. Translation-invariant operators
  • If $\lim\limits_{r\to 0}\ell(r)=\infty$, the operator is ultracontractive.
  • If $\lim\limits_{r\to 0}\ell(r)=c>0$, the operator is strongly pp–hypercontractive for every $p>1$, and also eventually ultracontractive.
@@
The first case was already established in [11, Corollary 4.5]. The second case follows from Theorem 5.1 and Section 6, or Theorem 6.6 for energy estimates and the optimal condition after which solutions get bounded.
Let us restrict our attention to the third case. If $\lim\limits_{r\to 0}\ell(r)=0$, it is easy to see that
@@ 7. Comments and possible extensions
Let us highlight what we regard as the most significant open question raised by this work: Are strong hypercontractivity and eventual ultracontractivity equivalent? Alternatively, can we find operators of the form (1.1) whose solutions to ($\textup{P}_{\mathcal{L}}$) enter every space $L^{q}$ for $q\in[1,\infty)$ in finite time, yet never reach $L^{\infty}$? According to our results, if such operators exist, we do not expect them to be translation-invariant. It may be necessary to impose some regularity conditions on the kernels in order to achieve $L^{\infty}$ bounds.
It is well established in the literature, under mild assumptions, that the standard instantaneous ultracontractivity and Nash-type inequalities are equivalent; see [13, 47, 29, 20]. As shown in [47, 20], a Nash-type inequality of the form
@@
As noted in the precedents Section 1.3, it was shown in [26] that the operator $\log(I-\Delta)$ satisfies a Nash-type inequality of the form
@@
where $A_{\rho},B_{\rho,N}>0$ depend on a constant $\rho>1$ and the dimension $N$. Observe that the left-hand side of (7.3) does not satisfy the integral condition (7.2), as expected. To the best of our knowledge, no connection has been established in the literature between eventual ultracontractivity and Nash inequalities of the form (7.3). It remains an open question whether such an equivalence exists and whether its underlying idea is related to the approach developed in this paper, namely, the use of a family of pp–Nash inequalities.
@@
A final and rather intriguing observation emerges when the problem is analyzed as a critical case of fractional Porous Media equations as
@@
where $m=m_{c}=\frac{(N-\sigma)_{+}}{N}$ is the critical exponent. Instantaneous ultracontractivity (known as $^{1}LL^{\infty}$ smoothing effect in this setting) appears when $m>m_{c}$, and there is lack of instantaneous regularization when $m\leq m_{c}$. What makes this particularly noteworthy is that for operators of zero or near-zero order, such as $\log(I-\Delta)$, we may think as $\sigma=0$, and consequently $m_{c}=1$. In this case, the linear scenario becomes critical.
@@
We can use this theorem to prove the existence of solutions to problem ($\textup{P}_{\mathcal{L}}$), where $H=L^{2}(\mathbb{R}^{N})$$D(\mathcal{L})=H_{\mathcal{L}}(\mathbb{R}^{N})$, and the operator $\mathcal{L}$ of type (1.1) is understood via $\langle\mathcal{L}f,g\rangle=\mathcal{E}(f,g)$ for $f,g\in H_{\mathcal{L}}(\mathbb{R}^{N})$. It is also worth mentioning that for this existence result to work, we do not require the kernel $J$ to be a Lévy kernel, it is enough that it be nonnegative and symmetric.
@@
Furthermore, if $p\geq 2$ and $u_{0}\in L^{\infty}(\mathbb{R}^{N})$, then $u\in\mathcal{C}([0,\infty);L^{p}(\mathbb{R}^{N}))$; if $1\leq p\leq 2$ and $u_{0}\in L^{1}(\mathbb{R}^{N})$, then $u\in\mathcal{C}([0,\infty);L^{p}(\mathbb{R}^{N}))$

### v2

@@ 6. Translation-invariant operators
  • (i)
    If $\lim\limits_{r\to 0}\ell(r)=\infty$, the operator is ultracontractive. Moreover, solutions belong to $\mathcal{C}^{\infty}(\mathbb{R}^{N})$ for any $t>0$.
  • (ii)
    If $\lim\limits_{r\to 0}\ell(r)=c>0$, the operator is strongly pp–hypercontractive for every $p>1$, and also eventually ultracontractive. Solutions improve in differentiability right after getting bounded.
  • (iii)
@@
(i) Ultracontractivity in this case was established in [11, Corollary 4.5]. As noticed there, the condition $\lim_{r\to 0}\ell(r)=\infty$ implies that $\ell$ is bounded from below, for $r$ sufficiently small, by a positive constant (which is a regularly varying function). Consequently, the estimate (6.9) applies, and we have
(6.16) $\lim_{|\xi|\to\infty}\frac{m(\xi)}{\log(|\xi|)}\geq c_{3}\lim_{r\to 0}\frac{\psi_{1}^{\ell}(r)}{\log(1/r)}=-c_{3}\lim_{r\to 0}r(\psi_{1}^{\ell})^{\prime}(r)=c_{3}\lim_{r\to 0}\ell(r)=\infty.$
Thus, the so-called Hartman-Wintner condition,
(6.17) $\lim_{|\xi|\to\infty}\frac{m(\xi)}{\log(|\xi|)}=\infty,$
holds, which es enough to guarantee that for any $t>0$ the fundamental solution is radial, nonincreasing in $r=|x|$ and bounded, and belongs to $C^{\infty}(\mathbb{R}^{N})$; see for instance [38].
(ii) Strong hypercontractivity follows in this case from Theorem 5.1 and Section 6. Theorem 6.6 also gives the strong hypercontractivity, as well as the optimal condition after which solutions get bounded and the differentiability improvement.
(iii) If $\lim\limits_{r\to 0}\ell(r)=0$, it is easy to see that
@@
There is a small subtlety in Theorem 6.7(iii): we show that the fundamental solution does not gain integrability, but we cannot a priori rule out this phenomenon for other solutions with $L^{p}$ initial data. To this end, we require an analogue of Section 6 valid in $L^{r}$, for arbitrary $r$, rather than only in $L^{\infty}$. In other words, we need to “invert” weak Young’s inequality.
Lemma 6.8 (Optimality of weak Young’s inequality).
Let $p,r\in\mathbb{R}$ such that $1\leq p<r<\infty$, and let $G\geq 0$ be a measurable function satisfying
(6.21) $\|G*f\|_{r}\leq C\|f\|_{p}$
for $C>0$ and any $f\in L^{p}(\mathbb{R}^{N})$. For $\lambda\geq 0$, define the sets
(6.22) $E_{\lambda}=\{x\in\mathbb{R}^{N}:G(x)>\lambda\}.$
Assume for any $\lambda\geq 0$ that
(6.23) $\mu(E_{\lambda}+E_{\lambda})\leq c\mu(E_{\lambda})$
for some $c>0$ independent of $\lambda$. Then $G\in L^{q,\infty}(\mathbb{R}^{N})$ for $q$ satisfying
(6.24) $1+\frac{1}{r}=\frac{1}{p}+\frac{1}{q}.$
Proof.
Let $f(x)=\chi_{E_{\lambda}}(x)$. From (6.21), we get that
(6.25) $\|G*\chi_{E_{\lambda}}\|_{r}\leq C\|\chi_{E_{\lambda}}\|_{p}=C\mu(E_{\lambda})^{1/p}.$
On the other hand,
(6.26) $G*\chi_{E_{\lambda}}(x)\geq\int_{E_{\lambda}}G(y)\chi_{E_{\lambda}}(x-y)\,dy\geq\lambda\big(\chi_{E_{\lambda}}*\chi_{E_{\lambda}}\big)(x).$
Now, notice that
(6.27) $\mu^{2}(E_{\lambda})=\|\chi_{E_{\lambda}}*\chi_{E_{\lambda}}\|_{1}=\int_{E_{\lambda}+E_{\lambda}}\chi_{E_{\lambda}}*\chi_{E_{\lambda}}\leq\|\chi_{E_{\lambda}}*\chi_{E_{\lambda}}\|_{r}\,\mu(E_{\lambda}+E_{\lambda})^{1-\frac{1}{r}}.$
Hence, using (6.23),
(6.28) $\|G*\chi_{E_{\lambda}}(x)\|_{r}\geq\lambda\,\frac{\mu^{2}(E_{\lambda})}{\mu(E_{\lambda}+E_{\lambda})^{1-\frac{1}{r}}}\geq\widetilde{c}\lambda\mu(E_{\lambda})^{1+\frac{1}{r}}.$
This, together with (6.25) implies that $\lambda\mu(E_{\lambda})^{1/q}$ is bounded independently of $\lambda\geq 0$, and thus we conclude the result. ∎
Hypothesis (6.23) is trivially satisfied if $E_{\lambda}$ is a convex set. In particular, if $\mathcal{H}_{t}$ is radial and nonincreasing in $r=|x|$ for each $t>0$, then the sets $\{x\in\mathbb{R}^{N}:H_{t}(x)>\lambda\}$ are balls and thus satisfy the condition. In order to ensure that $\mathcal{H}_{t}$ satisfies these requirements, we introduce the following definition, which allows us to state a sufficient condition.
Definition 6.9.
A measure $\mu(dx)$ on $\mathbb{R}^{N}$ is isotropic if there exists a function $f(r)$ $(0<r<\infty)$ such that $\mu(dx)=f(|x|)\,dx$ for $x\neq 0$. A measure $\mu(dx)$ on $\mathbb{R}^{N}$ is isotropic unimodal if $f(r)$ is nonincreasing ($\mu(\{0\}$ may be positive).
It turns out that the fundamental solution $\mathcal{H}_{t}$ is isotropic unimodal if and only if the kernel $\widetilde{J}(z)$ is isotropic unimodal; see [57]. Furthermore, $\widetilde{J}(z)$ is integrable if and only if $\int_{\{0\}}\mathcal{H}_{t}\,dx>0$ for $t>0$. With this in mind, we conclude the result below.
Corollary 6.10.
Let the kernel $\widetilde{J}$ defining the operator $\mathcal{L}$ be radial and nonincreasing in $r=|x|$, and assume that $\lim_{r\to\infty}\ell(r)=0$. Then,
  1. (i)
    If the kernel is integrable, solutions are only as regular as the initial data.
  2. (ii)
    If the kernel is not integrable, then solutions do not improve in integrability.
In any case, the operator is not pp–hypercontractive for any $p\geq 1$.
Proof.
(i) If the kernel is integrable, the solution to ($\textup{P}_{\mathcal{L}}$) is given by $u(t)=e^{-t}u_{0}+v(t)*u_{0}$ for some function $v$ (the regular part of the heat kernel); see for instance [14]. Therefore, $u(t)$ is only as good as the initial datum is.
(ii) If the kernel is not integrable, then we can guarantee that the fundamental solution $\mathcal{H}_{t}$ is radial and nonincreasing in $|x|$ for all $t>0$, since it has no Dirac delta at the origin. Thus, if solutions starting from initial data in $L^{p}$ with $p\geq 1$ improved in integrability, Section 6 would force $\mathcal{H}_{t}$ to belong to $L^{q}(\mathbb{R}^{N})$ for some $q>1$. However, this contradicts Theorem 6.7. ∎
Hence, not only is $\widetilde{J}(z)\sim|z|^{-N}$ for $|z|\ll 1$ the threshold for eventual ultracontractivity, but also for any kind of hypercontractivity. Observe that this previous result even works for p=1{p=1}, which is the most problematic case throughout the paper.
@@ 7. Comments and possible extensions
Let us highlight what we regard as the most significant open question raised by this work: Are strong hypercontractivity and eventual ultracontractivity equivalent? Alternatively, can we find operators of the form (1.1) whose solutions to ($\textup{P}_{\mathcal{L}}$) enter every space $L^{q}$ for $q\in[1,\infty)$ in finite time, yet never reach $L^{\infty}$? According to our results, if such operators exist, we do not expect them to be translation-invariant. The study of how regular solutions can become with only ellipticity assumptions in the spirit of Theorem 5.1 and Theorem 5.4, would correspond, in this setting, to the parabolic De Giorgi–Nash–Moser theory for local divergence-form operators. However, without additional regularity hypotheses on the kernel, at most $L^{\infty}$ is expected in our nonlocal framework, and even this may require some additional assumptions.
It is well established in the literature, under mild conditions, that the standard instantaneous ultracontractivity and Nash-type inequalities are equivalent; see [13, 53, 32, 20]. As shown in [53, 20], a Nash-type inequality of the form
@@
As noted in Section 1.3, it was shown in [29] that the operator $\log(I-\Delta)$ satisfies a Nash-type inequality of the form
@@
where $A_{\rho},B_{\rho,N}>0$ depend on a constant $\rho>1$ and the dimension $N$. Observe that the left-hand side of (7.3) does not satisfy the integral condition (7.2), as expected. To the best of our knowledge, no connection has been established in the literature between eventual ultracontractivity and Nash inequalities of the form (7.3). It remains open whether such an equivalence exists, and whether it is related to our approach based on a family of pp–Nash inequalities.
@@
A final and rather intriguing observation emerges when the problem is analysed as a critical case of fractional Porous Media equations as
@@
where $m=m_{c}=\frac{(N-\sigma)_{+}}{N}$ is the critical exponent. Instantaneous ultracontractivity (known as $^{1}LL^{\infty}$ smoothing effect in this setting) appears when $m>m_{c}$, and there is lack of instantaneous regularization when $m\leq m_{c}$. What makes this particularly noteworthy is that for operators of zero or $0^{+}$-order, such as $\log(I-\Delta)$, we may think of $\sigma$ as 0, and consequently $m_{c}=1$. In this case, the linear scenario becomes critical.
@@
We can use this theorem to prove the existence of solutions to problem ($\textup{P}_{\mathcal{L}}$), where $H=L^{2}(\mathbb{R}^{N})$$D(\mathcal{L})=H_{\mathcal{L}}(\mathbb{R}^{N})$, and the operator $\mathcal{L}$ of type (1.1) is understood via $\langle\mathcal{L}f,g\rangle\coloneq\mathcal{E}(f,g)$ for $f,g\in H_{\mathcal{L}}(\mathbb{R}^{N})$. It is also worth mentioning that for this existence result to work, we do not require the kernel $J$ to be a Lévy kernel, it is enough that it be nonnegative and symmetric.
@@
Furthermore, if $p\geq 2$ and $u_{0}\in L^{\infty}(\mathbb{R}^{N})$, then $u\in\mathcal{C}([0,\infty);L^{p}(\mathbb{R}^{N}))$; if $1\leq p\leq 2$ and $u_{0}\in L^{1}(\mathbb{R}^{N})$, then $u\in\mathcal{C}([0,\infty);L^{p}(\mathbb{R}^{N}))$.
@@ Acknowledgements
J. Ruiz-Cases would also like to thank X. Ros-Oton for the fruitful discussions related to Section 6 during a visit to Universidad de Barcelona funded by RED2024-153842-T.